OpenStudy (anonymous):

i hate throwing up at morning lol but have a qs to do cud anyone help ? will fan+medal

OpenStudy (anonymous):

Verify that (√2 − i) − i(1 −√2i) = −2i

OpenStudy (owlcoffee):

Okay let's begin this: \[(\sqrt{2}-i)-i(1-\sqrt{2}i)\] I'll begin this by applying distributive property on the second member with that "i": \[(\sqrt{2}-i)-i+\sqrt{2}i ^{2}\] Now, we'll have to take back what we learned in complet numbers, we know that: \[i=\sqrt{-1}\] But what will happen if I square it, if i=sqrt (-1) then: \[i ^{2}=(\sqrt{-1})^{2}\] \[i ^{2}=-1\] Let's go back to the problem and apply what we just deduced: \[(\sqrt{2}-i)-i+\sqrt{2}(-1)\] Now it's just a matter of operating: \[(\sqrt{2}-i)-i-\sqrt{2}\] \[\sqrt{2}-i-i-\sqrt{2}=-i-i=2i\] Therefore: \[(\sqrt{2}-i)-i(1-\sqrt{2}i)=2i\]

OpenStudy (anonymous):

sweet ! i got it nw ty alote ^^